OpenStudy (anonymous):

Could someone please help me with this question? 5.00 g of nitrogen is completely converted into an oxide of nitrogen. The mass of the oxide formed is 19.3g. What is the empirical formula of the oxide?

OpenStudy (chmvijay):

LOL its N2O5

OpenStudy (chmvijay):

its like trial and error method just check for the formaion of different oxides oxide N2 + O2 ------> 2 NO in the above reaction 28 gram of N2 gives rise too 60 gram of NO then if u burn 5 gram of N2 = 60*5 / 28 =10.71 gram of NO then i checked for N2+O---->N2O in the above reaction 28 gram of N2 gives rise too 44 gram of N2O then if u burn 5 gram of N2 = 44*5 / 28 = 7.8 gram of N2O then i tried for the following and got ur result N2 +O2------> N2O5 in the above reaction 28 gram of N2 gives rise too 108 gram of N2O5( nitrogen petoxide) then if u burn 5 gram of N2 = 108*5 / 28 =19.28 gram of N2O5 the answer is N2O5

OpenStudy (mos1635):

\[N _{2} +\frac{ x }{ 2 } O _{2} \rightarrow N _{2}O _{x}\]

OpenStudy (mos1635):

1 mole of N2 gives 1 mole of oxide there for n1=n2 m1/Mr1 = m2/Mr2 5/28 =19.3/(28+16x)

OpenStudy (mos1635):

\[x \approx5\]

OpenStudy (chmvijay):

LOL nice :)